Calculate the Rms value of a sinusoidal current of maximum value of 20A


1
Expert's answer
2021-06-16T04:17:01-0400
"I_{rms}=\frac{I}{\sqrt{2}}\\I_{rms}=\frac{20}{\sqrt{2}}=14.14\ A"